OpenStudy (anonymous):

Is 5/13 greater than, less than or equal to 1/4???

OpenStudy (anonymous):

to know without using a calculator, make their denominator equal. find the lowest common multiple of 13 and 4. turns our its 13x4 = 52 so we want to put them both over 52. notice that im multiplying by ONE technically. \[\frac{ 5 }{ 13 } = \frac{ 5 }{ 13 } \times \frac{ 4 }{ 4 } = \frac{ 20 }{ 52 }\]and\[\frac{ 1 }{ 4 } = \frac{ 1 }{ 4 } \times \frac{ 13 }{ 13 } = \frac{ 13 }{ 52 }\] which is bigger? :)

OpenStudy (anonymous):

Thank you